OpenStudy (bimbim2697):

FIND THE THIRD,FIFTH AND TENTH TERM SEQUENCE DESCRIBED BY EACH RULE A(n)=4+(N-1)(-5) A(N)=-2(n-1)(-5)

OpenStudy (hoblos):

for the first one: A(3)= 4 + (2)(-5) = -6 A(5)= 4 + 4(-5) = -16 A(10)= -41

OpenStudy (hoblos):

why dont you try the second and tell me what you get lets see if you've got the way

OpenStudy (bimbim2697):

A(3)=-2(4)(-5)=40

OpenStudy (bimbim2697):

AM I RIGHT?

OpenStudy (hoblos):

n-1 = 3-1 = 4??!!!!

OpenStudy (bimbim2697):

HUH?

OpenStudy (bimbim2697):

OH CRAP I GOT IT ...

OpenStudy (bimbim2697):

A(3)=-2(2)(-5)=20

OpenStudy (hoblos):

yup now it is correct :D

OpenStudy (bimbim2697):

ARE U A TESCHER HOBLOS?

OpenStudy (hoblos):

!!!nope

OpenStudy (bimbim2697):

YOU TEACHER WAY!!!!!!!!!!!!!!!!!!!!!!!!!1 BETTER THAN MY TEACHER

OpenStudy (hoblos):

hahahaha thanks :D

OpenStudy (bimbim2697):

A(5)=-2(4)(-5)=40

OpenStudy (hoblos):

yup.... one more

OpenStudy (bimbim2697):

A(10)=-2(9)(-5)=90

OpenStudy (hoblos):

good job :D

OpenStudy (bimbim2697):

:p

OpenStudy (bimbim2697):

I HAVE A FEW MORE BUT DIFFERENT STYLE HAHA

OpenStudy (hoblos):

post them and i'll see if i can help ;)